Two variables P and Q has values 30±2 and 10.0 respectively, calculate the percentage error in the following (i) PQ (ii) P2 (iii)
1
Expert's answer
2020-10-05T10:57:00-0400
"PQ=(30\pm2)10=300\pm20,\\\space\\
\epsilon=\frac{20}{300}\cdot100\%=6\%."

"P^2=100,\\
\epsilon=\frac{0}{100}\cdot100\%=0."